Skip to Search

 
diagonal spacer

Information Technology Staff Meeting Notes
November 16, 2009

Present: Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Bill Covey, Matthew Daley, Debra Fetzer, Winston Harris, Michael Jay, Bobbie Meng, Chris Nicolich, Amy Polk, Josh Spurgin, Evan Wack, Stephen Williams
Announcements
  1. Move updates - discussion of the latest floor plan, attached to Meeting Agenda handout. We will continue to get status reports as things progress.
  2. Trish will be back tomorrow.
  3. We assume Michael Howell will be back this week.
  4. Evan will be out for several weeks in the near future when his wife has the baby.
  5. Our new hardware technician, Josh, started today.
  6. New programmer, Shruti, has been hired and starts Monday.
  7. We need to think about the IT Dept holiday party.
Ongoing IT Updates

Orders
  1. Order for 22 workstations with dual monitors has been approved and is in process.
  2. Hardware for lobby installation - 2 monitors have been received and are stored in the vault. We are still waiting to receive the workstation. Also waiting for network drops to be pulled. When everything is in place, someone from the Hardware Unit will be responsible for the installation. Since there were inquiries as to when the equipment will arrive, it is possible they will want this set up for the dedication ceremony of the Smathers Library Lobby Renovations on December 2.
  3. Wish list items - some items (scanners, etc.) are processing via the IT budget, not FCLA.
  4. MSL order to expand the MSL SAN to match the current - on hold pending the estimate for the virtualization all of our servers at CNS. New quotes will be necessary if we end up going ahead with the expansion order. We may have the option to co-locate our servers in the CSE building. The problem with CSE is the close proximity to the back-up server. Chuck Frasier and Judy met; he said we could have space in the East campus location. All yet to be determined.
Hardware
  1. Public printing/copying - Canon says the fix sent by Pharos has script errors. It was sent back to Canon on November 2nd. They are supposed to get back to us today with a projected time for a fix - probably in December.
  2. West printing issue - log-on script error - hang time on black/white printer: it could not be found. Seems to be a log-in delay. Could be an issue with the group policy setting. Will log a ticket in Grover for Amy.
  3. Local DHCP in East - Logan has been busy working on DLC things - he will work on this with Amy.
  4. Video conferencing - it's off Evan's plate - the proposal has gone out. Ticket generated for Facilities to get the power going.
  5. Win7 - laptop image - will start on it tomorrow.
  6. Workstation Support Unit has been tasked with sanitizing all public workstations. Josh notes that they use Sanigard at SFC - it sprays on and protects for two weeks. We will look into getting that product.
  7. Icon for value add-on - no word from Rich.
  8. Backup Exec - running out of space for backup to disk - Logan looking into it. We are using 20 tapes. There is corruption on the config management folders. We are trying to migrate that data to DVD.
  9. Matthew has a ticket where students asked for Chrome to be loaded on the public machines. We will add it as the public machines are re-imaged.
Software
  1. Grover Purchasing mods are done. Chris will demo tomorrow to Fiscal Services. Sonia seemed quite pleased with it - just had a few changes.
  2. Grover Facilities mod - we have received Rob's wish list. This project will take longer as we will have to transfer the data from their current ticketing system.
  3. Grover Acquisitions mod - no ticket generated; details unknown.
  4. Computer moves in West - CAT needs to be updated. Please pass the details on to Winston and he will update it.
  5. Will cautions to make sure all requests, etc. are entered in Grover. Mark got pulled into a grant proposal which required server support.
  6. VIVO - Stephen returned from Indiana and is now on his way to St. Louis. Logan is registering vivo.ufl.edu with CNS today.
  7. Macros - ongoing. Per Winston, all should be done by the end of the month.
  8. Room reservation app - pending - sent to Stacey Ewing.
  9. JABBR - need a production Linux box as an app server. Logan to get it done this week.
  10. Issues with the Incident Report Form - it seems there are multiple versions of the form out there. We killed one.
  11. Maps Library - Toppo program - it is only usable on campus and they want it to work for off-campus folks.
Network Server Admin
  1. Meeting with CNS on Wednesday to discuss virtualization cost.
Web Projects
  1. Feedback is slow on outdated info. The deadline given: December 1st. Matthew to inform Will on departments that have not complied. Permissions are a problem when Matthew is asked to delete data.
  2. Patron Grover - several outstanding tickets. People are unsure how to respond. Laborious and confusing.
  3. Matthew has been asked to generate ARL stats for Steve Carrico - how many visits to the Library home page. Need to run a click track.
  4. Content management system - nothing yet. Need to work on a template.
  5. Intranet - no word.
DLC General Issues
  1. FTP - Logan will talk to CNS this week.
  2. JPG2K - this week.
  3. Streaming video - pending outside forces - awaiting the Dean's decision.
IT Security Management
  1. Meeting held.

 

University of Florida Home Page